OpenStudy (anonymous):

Give the center radius of each equation.Then sketch the graph. (x-3)square + (y+4)square minus 9=0

OpenStudy (anonymous):

\[(x-3)^2+(y+4)^2=9\] if center = (h,k) and radius = r, then the general form is \[(x-h)^2+(y-k)^2=r^2\] Thus center=(3,-4) and r=3

OpenStudy (anonymous):

why is 4 negative and 3 positive and thank so much

OpenStudy (anonymous):

because it must be in the form (x-h)^2 and so for (x-h) to equal (x-3), h must be positive 3. Similarly, if (y+4)=(y-k), then k=-4

OpenStudy (anonymous):

ok how is this supose to be graphed because im looking at the example of the graph and it very difficult

OpenStudy (anonymous):

do you have a compass?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well, it's going to be a circle, so I would, starting from (3,-4), go 3 units up, down, left and right.

OpenStudy (anonymous):

then draw a curve through those four points by hand and make it as circular as you can?

OpenStudy (anonymous):

ok one more question same problem different equation

OpenStudy (anonymous):

x square + y square = 36

OpenStudy (anonymous):

ok, this one is actually a little simpler. See, to get it in the standard form, all we do is rewrite it as \[(x-0)^2+(y-0)^2=6^2\] now clearly it's the same equation as the one you gave me, just a little more convenient

OpenStudy (anonymous):

and from this we determine the center is (0,0) and the radius is 6

OpenStudy (anonymous):

why is it 6 when you square root it

OpenStudy (anonymous):

sorry dumb question

OpenStudy (anonymous):

i see lol

OpenStudy (anonymous):

that's alright, as long as you understand it now

OpenStudy (anonymous):

and if you don't, i'd be happy to answer anything else

OpenStudy (anonymous):

man you have been a big help ill be sure to be back on this mabey in a hour or so or minutes thanks :)